# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 3.1 Conjugacy properties of sue distributions in multivariate linear models
Let us first study the conjugacy properties of sue distributions under general multivariate linear models of the form
@@
where ${\bf{y}}=(y_{1},\ldots,y_{n})^{\top}\in\mathbb{R}^{n}$ is the response vector, ${\bf{X}}\in\mathbb{R}^{n\times p}$ corresponds to a known design matrix, ${\bm{\beta}}\in\mathbb{R}^{p}$ denotes a vector of unknown parameters, often referred to as the regression coefficients, and ${\bm{\varepsilon}}\in\mathbb{R}^{n}$ is the error vector. Current results in Bayesian inference under the above model have established conjugacy of Gaussian or sun priors for ${\bm{\beta}}$ when combined with Gaussian or sun noise vectors ${\bm{\varepsilon}}$ [4]. Although these advancements cover a broad range of models, in practice, it is of interest to consider alternative representations within the wider elliptical or skew–elliptical family, which account for heavier tails and ensure increased robustness. However, conjugacy remains unexplored in these larger classes, undermining advancements in tractable Bayesian inference. Proposition 1 below covers such a gap.
@@
Then, when ${\bf{y}}$ is defined as in Equation (20), it follows that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with
@@
Before proving Proposition 1, it shall be emphasized that the above results, along with those provided in Propositions 2 and 3, are purposely stated in a highly–general form in order to derive a comprehensive conjugacy theory for sue distributions that is of broader and independent interest in expanding the theoretical analysis of such a family. As clarified in Examples 16, priors and likelihoods of potential interest in practice are only a subset of the general results in Propositions 13. More specifically, setting $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$, $\bm{\tau}=\bm{0}$, ${\bm{\xi}}_{{\bm{\varepsilon}}}=\bm{0}$, and either ${\bm{\Delta}}_{\bm{\beta}}=\bm{0}$ or ${\bm{\Delta}}_{\bm{\varepsilon}}=\bm{0}$, would be sufficient to recover most of the priors and likelihoods of direct interest in applications.
@@
To prove Equation (22) in Proposition 1, first notice that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}={\bf{A}}({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$, where ${\bf{A}}$ denotes a known matrix of dimension $(p+n)\times(p+n)$ with blocks ${\bf{A}}_{11}={\bf{I}}_{p}$, ${\bf{A}}_{12}=\bm{0}$, ${\bf{A}}_{21}={\bf{X}}$ and ${\bf{A}}_{22}={\bf{I}}_{n}$. Combining such a representation with the closure under linear combination properties of sue distributions presented in Lemma 1, we have that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bf{A}}{\bm{\xi}},{\bf{A}}\bm{\Omega}{\bf{A}}^{\top},{\bm{\Delta}}_{\bf{A}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ where ${\bf{A}}{\bm{\xi}}={\bm{\xi}}^{\dagger}$, ${\bf{A}}\bm{\Omega}{\bf{A}}^{\top}=\bm{\Omega}^{\dagger}$ and ${\bm{\Delta}}_{\bf{A}}={\bm{\Delta}}^{\dagger}$. As a result, the prior distribution for ${\bm{\beta}}$ in Proposition 1 follows directly from the closure under marginalization of the sue family outlined in Lemma 1. Similarly, the likelihood $({\bf{y}}\mid{\bm{\beta}})$ and posterior $({\bm{\beta}}\mid{\bf{y}})$ in Proposition 1 can be readily derived by applying the closure under conditioning properties in Lemma 2 to the joint sue distribution for $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}$ presented in Proposition 1, with parameters ${\bm{\xi}}^{\dagger}$, $\bm{\Omega}^{\dagger}$ and ${\bm{\Delta}}^{\dagger}$ partitioned as in (22). ∎
@@
As anticipated in Section 3, the results in Proposition 1 clarify that the joint distribution for ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ requires some form of dependence to guarantee conjugacy. In this respect, notice that even when ${\bm{\xi}}=\bm{0}$, $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$, by the closure under conditioning properties of sue, we have $({\bm{\varepsilon}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}(\bm{0},\bm{\Omega}_{\bm{\varepsilon}},\bm{0},\bm{\tau},\bar{\bm{\Gamma}},g_{Q_{\bm{\beta}}({\bm{\beta}})}^{(n+q)})$, which clarifies that a weak form of dependence persists in the conditional density generator. Nonetheless, such a form of higher–level dependence still allows to include within the results in Proposition 1 interesting statistical models based on uncorrelated ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ vectors. Recalling the expression for the covariance matrix of sue distributions in Equation (7), a sufficient condition to retrieve these uncorrelated representations is to consider $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$ and assume either ${\bm{\Delta}}_{\bm{\beta}}=\bm{0}$ or ${\bm{\Delta}}_{\bm{\varepsilon}}=\bm{0}$. When both ${\bm{\Delta}}_{\bm{\beta}}$ and ${\bm{\Delta}}_{\bm{\varepsilon}}$ are $\bm{0}$, and also $\bm{\tau}=\bm{0}$, by point (i) in Lemma 5, $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ reduces to an elliptical distribution. As such, conjugacy under this latter class can be obtained as a special case of Proposition 1.
@@
The above discussion clarifies that full independence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ cannot be generally enforced if the objective is to obtain broad conjugacy results as in Proposition 1 that hold for the whole sue family. Nonetheless, in the specific setting of Gaussian density generators, which leads to the sub–class of sun distributions, such a full independence can be enforced without undermining conjugacy. As discussed in Section 2.2.1, under this specific choice, the conditional density generator coincides with the unconditional one, thus allowing to enforce independence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ while preserving conjugacy. This is clear from the results in Anceschi et al. [4], that establish sun conjugacy via a classical Bayes rule perspective, without requiring to specify a joint distribution for ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ or, alternatively, ${\bm{\beta}}$ and ${\bf{y}}$. As illustrated in Example 1 below, these conjugacy results can be obtained as a particular case of those in Proposition 1.
Example 1 (Multivariate unified skew–normal (sun) conjugacy).
@@
The Supplementary Materials of Anceschi et al. [4] present an example based on a classical linear regression with skew–normal errors, i.e., $(y_{i}\mid{\bm{\beta}})\sim{\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$, independently for $i\in\{1,\ldots,n\}$, and, consistent with our notation, $\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$ prior for ${\bm{\beta}}$. This model yields a likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})\propto\phi_{n}({\bf{y}}-{\bf{X}}{\bm{\beta}};\sigma^{2}{\bf{I}}_{n})\,\Phi_{n}(\alpha{\bf{y}}-\alpha{\bf{X}}{\bm{\beta}};\sigma^{2}{\bf{I}}_{n})$, proportional to a $\mbox{SUN}_{n,n}({\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},\alpha\sigma{\bf{I}}_{n},\bm{0},(1+\alpha^{2})\sigma^{2}{\bf{I}}_{n})$ density. Leveraging Lemma 4, such a sun is equivalent to $({\bf{y}}\mid{\bm{\beta}})\sim\mbox{SUN}_{n,n}({\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},[\alpha/(1+\alpha^{2})^{1/2}]{\bf{I}}_{n},\bm{0},{\bf{I}}_{n})$. Before showing that this Bayesian formulation is a special case of the broader family of models and priors in Proposition 1, it shall be emphasized that this construction also comprises classical multivariate Gaussian priors for ${\bm{\beta}}$, when ${\bm{\Delta}}_{{\bm{\beta}}}=\bm{0}$, and Gaussian linear regression for ${\bf{y}}$ if $\alpha=0$. Replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix also leads to general multivariate versions of such models. This yields an important class of routinely–implemented formulations.
@@
where $\bar{\alpha}=\alpha/(1+\alpha^{2})^{1/2}$. Adapting Proposition 1 to this setting, yields $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim\smash{{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}})},$ with
@@
As a direct consequence of the closure under linear combinations of sue, and hence sun, distributions, together with point (ii) in Lemma 5, the above formulation implies
@@
where the marginal for ${\bm{\beta}}$ coincides with the prior considered in Anceschi et al. [4], whereas ${\bm{\varepsilon}}$ corresponds to a noise vector comprising $n$ independent skew–normals. Similarly, by applying to the above parameters the expressions for those of $({\bf{y}}\mid{\bm{\beta}})$ in Proposition 1, and recalling Remark 2, yields after standard calculations the following likelihood
@@
Such a likelihood is proportional to the one considered in Anceschi et al. [4] for the general skew–normal regression setting, which includes the Gaussian as a special case and can be readily extended to more general multivariate models. As such, Proposition 1 also covers sun conjugacy properties in commonly–implemented linear models.
Example 2 (Multivariate unified skew–$t$ (sut) conjugacy).
As stated in Corollary 1, by specializing Proposition 1 to the sut sub–family presented in Section 2.2.2, it is possible to derive novel conjugacy results not yet explored in the current literature, along with specific examples of potential interest in applications.
@@
Consider the linear regression model defined as in (20), and let $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUT}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu),$ with parameters ${\bm{\xi}},\bm{\Omega},{\bm{\Delta}}$ partitioned as in (21). Then, the induced prior distribution is ${\bm{\beta}}\sim{{\mbox{SUT}}}_{p,q}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},{\bm{\Delta}}_{\bm{\beta}},\bm{\tau},\bar{\bm{\Gamma}},\nu)$, the likelihood corresponds to $({\bf{y}}\mid{\bm{\beta}})\sim{{\mbox{SUT}}}_{n,q}({\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\alpha_{\bm{\beta}}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\alpha_{\bm{\beta}}^{-1/2}\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},{\bar{\bm{\Gamma}}}_{{\bf{y}}\mid{\bm{\beta}}},\nu+p)$, whereas the resulting posterior distribution coincides with $({\bm{\beta}}\mid{\bf{y}})\sim{{\mbox{SUT}}}_{p,q}({\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}},\alpha_{\bf{y}}\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}},{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}},\alpha_{\bf{y}}^{-1/2}\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}},{\bar{\bm{\Gamma}}}_{{\bm{\beta}}\mid{\bf{y}}},\nu+n)$. In these expressions, $\alpha_{\bm{\beta}}=[\nu+Q_{\bm{\beta}}({\bm{\beta}})]/(\nu+p)$ and $\alpha_{\bf{y}}=[\nu+Q_{\bf{y}}({\bf{y}})]/(\nu+n)$, with $Q_{\bm{\beta}}({\bm{\beta}})=({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})^{\top}\bm{\Omega}_{\bm{\beta}}^{-1}({\bm{\beta}}-{\bm{\xi}}_{\bm{\beta}})$ and $Q_{\bf{y}}({\bf{y}})=({\bf{y}}-{\bf{X}}{\bm{\xi}}_{\bm{\beta}}-{\bm{\xi}}_{\bm{\varepsilon}})^{\top}\bm{\Omega}_{\bf{y}}^{-1}({\bf{y}}-{\bf{X}}{\bm{\xi}}_{\bm{\beta}}-{\bm{\xi}}_{\bm{\varepsilon}})$. The remaining parameters are defined as in Proposition 1.
@@
The proofs follow directly by replacing the generic density generators in Proposition 1 with those of the Student’s $t$ distribution presented in Section 2.2.2. Alternatively, it is possible to prove the statement leveraging the specific properties of unified skew–$t$ distributions in Proposition 11 of Wang et al. [62]. ∎
Corollary 1 states a general conjugacy result which further includes those of suns as a limiting case, provided that sut distributions converge to suns when $\nu\to\infty$ [e.g., 62]. In addition, suitable constraints on the parameters of the joint sut distribution for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ in Corollary 1, yield priors and likelihoods of potential practical interest. In particular, consider $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUT}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu),$ with
@@
${\bm{\beta}}\sim\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu),\qquad{\bm{\varepsilon}}\sim{{\mbox{SUT}}}_{n,q}(\bm{0},\bm{\Omega}_{{\bm{\varepsilon}}},{\bm{\Delta}}_{{\bm{\varepsilon}}},\bm{0},{\bar{\bm{\Gamma}}}_{{\bm{\varepsilon}}},\nu),$
which yield a Bayesian multivariate regression model ${\bf{y}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}}$ with Student’s $t$ prior on ${\bm{\beta}}$ and unified skew–$t$ residuals in ${\bm{\varepsilon}}$, uncorrelated with ${\bm{\beta}}$. In the above expression, $\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu)$ denotes the $p$–variate Students’ $t$ distribution with location ${\bm{\xi}}_{\bm{\beta}}$, scale $\bm{\Omega}_{\bm{\beta}}$ and degrees of freedom $\nu$. By Corollary 1, this implies the sut likelihood

### v3

@@ 3.1 Conjugacy properties of SUE distributions in multivariate linear models
Let us first study the SUE conjugacy properties under general multivariate linear models of the form
@@
where ${\bf{y}}=(y_{1},\ldots,y_{n})^{\top}\in\mathbb{R}^{n}$ is the response vector, ${\bf{X}}\in\mathbb{R}^{n\times p}$ corresponds to a known design matrix, ${\bm{\beta}}\in\mathbb{R}^{p}$ denotes a vector of unknown parameters, often referred to as the regression coefficients, and ${\bm{\varepsilon}}\in\mathbb{R}^{n}$ is the error vector. Current results in Bayesian inference under the above model have established conjugacy of Gaussian or SUN priors for ${\bm{\beta}}$ when combined with Gaussian or SUN noise vectors ${\bm{\varepsilon}}$ [4]. Although these advancements cover a broad range of models, in practice, it is of interest to consider alternative representations within the wider elliptical or skew-elliptical family, which account for heavier tails and ensure increased robustness. However, conjugacy remains unexplored in these larger classes, undermining advancements in tractable Bayesian inference. Proposition 1 below covers such a gap.
@@
Then, when ${\bf{y}}$ is defined as in (20), it follows that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ with
@@
Before proving Proposition 1, it shall be emphasized that the above results, along with those provided in Propositions 2 and 3, are purposely stated in a highly-general form in order to derive a comprehensive conjugacy theory for SUE distributions that is of broader and independent interest in expanding the theoretical analysis of such a family. As clarified in Examples 16, priors and likelihoods of potential interest in practice are only a subset of the general results in Propositions 13. More specifically, setting $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$, $\bm{\tau}=\bm{0}$, ${\bm{\xi}}_{{\bm{\varepsilon}}}=\bm{0}$, and either ${\bm{\Delta}}_{\bm{\beta}}=\bm{0}$ or ${\bm{\Delta}}_{\bm{\varepsilon}}=\bm{0}$, would be sufficient to recover most of the priors and likelihoods of direct interest in applications.
@@
To prove (22) in Proposition 1, first notice that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}={\bf{A}}({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$, where ${\bf{A}}$ denotes a known matrix of dimension $(p+n)\times(p+n)$ with blocks ${\bf{A}}_{11}={\bf{I}}_{p}$, ${\bf{A}}_{12}=\bm{0}$, ${\bf{A}}_{21}={\bf{X}}$ and ${\bf{A}}_{22}={\bf{I}}_{n}$. Combining such a representation with the closure under linear combination properties of SUE distributions presented in Lemma 1, we have that $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim{{\mbox{SUE}}}_{p+n,q}({\bf{A}}{\bm{\xi}},{\bf{A}}\bm{\Omega}{\bf{A}}^{\top},{\bm{\Delta}}_{\bf{A}},\bm{\tau},\bar{\bm{\Gamma}},g^{(p+n+q)}),$ where ${\bf{A}}{\bm{\xi}}={\bm{\xi}}^{\dagger}$, ${\bf{A}}\bm{\Omega}{\bf{A}}^{\top}=\bm{\Omega}^{\dagger}$ and ${\bm{\Delta}}_{\bf{A}}={\bm{\Delta}}^{\dagger}$. As a result, the prior distribution for ${\bm{\beta}}$ in Proposition 1 follows directly from the closure under marginalization of the SUE family outlined in Lemma 1. Similarly, the likelihood $({\bf{y}}\mid{\bm{\beta}})$ and posterior $({\bm{\beta}}\mid{\bf{y}})$ in Proposition 1 can be readily derived by applying the closure under conditioning properties in Lemma 2 to the joint SUE distribution for $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}$ presented in Proposition 1, with parameters ${\bm{\xi}}^{\dagger}$, $\bm{\Omega}^{\dagger}$ and ${\bm{\Delta}}^{\dagger}$ partitioned as in (22). ∎
@@
As anticipated within Section 3, the results in Proposition 1 clarify that the joint distribution for ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ requires some form of dependence to guarantee conjugacy. In this respect, notice that even when ${\bm{\xi}}=\bm{0}$, $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$ and ${\bm{\Delta}}=\bm{0}$, by the closure under conditioning properties of the unified skew-elliptical family, it follows that $({\bm{\varepsilon}}\mid{\bm{\beta}})\sim{{\mbox{SUE}}}_{n,q}(\bm{0},\bm{\Omega}_{\bm{\varepsilon}},\bm{0},\bm{\tau},\bar{\bm{\Gamma}},g_{Q_{\bm{\beta}}({\bm{\beta}})}^{(n+q)})$, which clarifies that a weak form of dependence persists in the conditional density generator. Nonetheless, such a form of higher-level dependence still allows to include within the results in Proposition 1 interesting models with uncorrelated ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ vectors. Recalling the expression for the SUE covariance matrix in (7), a sufficient condition to retrieve these uncorrelated representations is to assume either ${\bm{\Delta}}_{\bm{\beta}}=\bm{0}$ or ${\bm{\Delta}}_{\bm{\varepsilon}}=\bm{0}$, and set $\bm{\Omega}_{{\bm{\varepsilon}}{\bm{\beta}}}=\bm{\Omega}^{\top}_{{\bm{\beta}}{\bm{\varepsilon}}}=\bm{0}$. When both ${\bm{\Delta}}_{\bm{\beta}}$ and ${\bm{\Delta}}_{\bm{\varepsilon}}$ are $\bm{0}$, and also $\bm{\tau}=\bm{0}$, by point (i) in Lemma 5, $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ reduces to an elliptical distribution. As such, conjugacy under this latter class can be obtained as a special case of Proposition 1.
@@
The above discussion clarifies that full independence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ cannot be generally enforced if the objective is to obtain broad conjugacy results as in Proposition 1 that hold for the whole SUE family. Nonetheless, in the specific setting of Gaussian density generators, which leads to the sub-class of SUN distributions, such a full independence can be enforced without undermining conjugacy. As discussed in Section 2.2.1, under this specific choice, the conditional density generator coincides with the unconditional one, thus allowing to enforce independence between ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ while preserving conjugacy. This is clear from the results in Anceschi et al. [4], that establish SUN conjugacy via a classical Bayes rule perspective, without requiring to specify a joint distribution for ${\bm{\beta}}$ and ${\bm{\varepsilon}}$ or, alternatively, ${\bm{\beta}}$ and ${\bf{y}}$. As illustrated in Example 1 below, these conjugacy results can be obtained as a particular case of those in Proposition 1.
Example 1 (Multivariate unified skew-normal (SUN) conjugacy).
@@
The supplementary materials of Anceschi et al. [4] present an example based on a classical linear regression with skew-normal errors, i.e., $(y_{i}\mid{\bm{\beta}})\sim{\mbox{SN}}({\bm{x}}_{i}^{\top}{\bm{\beta}},\sigma^{2},\alpha)$, independently for $i\in\{1,\ldots,n\}$, and, consistent with our notation, $\mbox{SUN}_{p,q}({\bm{\xi}}_{{\bm{\beta}}},\bm{\Omega}_{{\bm{\beta}}},{\bm{\Delta}}_{{\bm{\beta}}},\bm{\tau}_{{\bm{\beta}}},\bar{\bm{\Gamma}}_{{\bm{\beta}}})$ prior for ${\bm{\beta}}$. This model yields a likelihood $\mathcal{p}({\bf{y}}\mid{\bm{\beta}})\propto\phi_{n}({\bf{y}}-{\bf{X}}{\bm{\beta}};\sigma^{2}{\bf{I}}_{n})\,\Phi_{n}(\alpha{\bf{y}}-\alpha{\bf{X}}{\bm{\beta}};\sigma^{2}{\bf{I}}_{n})$, proportional to a $\mbox{SUN}_{n,n}({\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},\alpha\sigma{\bf{I}}_{n},\bm{0},(1+\alpha^{2})\sigma^{2}{\bf{I}}_{n})$ density. Leveraging Lemma 4, such a SUN is equivalent to $({\bf{y}}\mid{\bm{\beta}})\sim\mbox{SUN}_{n,n}({\bf{X}}{\bm{\beta}},\sigma^{2}{\bf{I}}_{n},[\alpha/(1+\alpha^{2})^{1/2}]{\bf{I}}_{n},\bm{0},{\bf{I}}_{n})$. Before showing that this Bayesian formulation is a special case of the broader family of models and priors in Proposition 1, it shall be emphasized that this construction also comprises classical multivariate Gaussian priors for ${\bm{\beta}}$, when ${\bm{\Delta}}_{{\bm{\beta}}}=\bm{0}$, and Gaussian linear regression for ${\bf{y}}$ if $\alpha=0$. Replacing $\sigma^{2}{\bf{I}}_{n}$ with a full covariance matrix also leads to general multivariate versions of such models. This yields an important class of routinely-implemented formulations.
@@
where $\bar{\alpha}=\alpha/(1+\alpha^{2})^{1/2}$. Adapting Proposition 1 to this setting, yields $({\bm{\beta}}^{\top},{\bf{y}}^{\top})^{\top}\sim{{{\mbox{SUN}}}_{p+n,q+n}({\bm{\xi}}^{\dagger},\bm{\Omega}^{\dagger},{\bm{\Delta}}^{\dagger},\bm{\tau},\bar{\bm{\Gamma}})},$ with
@@
As a direct consequence of the closure under linear combinations of SUE, and hence SUN, distributions, together with point (ii) in Lemma 5 (see also Remark 2), the above formulation implies
@@
where the marginal for ${\bm{\beta}}$ coincides with the prior considered in Anceschi et al. [4], whereas ${\bm{\varepsilon}}$ corresponds to a noise vector comprising $n$ independent skew-normals. Similarly, by applying to the above parameters the expressions for those of $({\bf{y}}\mid{\bm{\beta}})$ in Proposition 1, and recalling Remark 2, yields after standard calculations the following likelihood
@@
Such a likelihood is proportional to the one considered in Anceschi et al. [4] for the general skew-normal regression setting, which includes the Gaussian as a special case and can be readily extended to more general multivariate models. As such, Proposition 1 also covers SUN conjugacy properties in commonly-implemented linear models.
Example 2 (Multivariate unified skew-$t$ (SUT) conjugacy).
As stated in Corollary 1, by specializing Proposition 1 to the SUT sub-family presented in Section 2.2.2, it is possible to derive novel conjugacy results not yet explored in the current literature, along with specific examples of potential interest in applications.
@@
Consider the model in (20), and assume $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUT}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu),$ with ${\bm{\xi}},\bm{\Omega},{\bm{\Delta}}$ partitioned as in (21). Then, the prior distribution for ${\bm{\beta}}$ is ${\bm{\beta}}\sim{{\mbox{SUT}}}_{p,q}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},{\bm{\Delta}}_{\bm{\beta}},\bm{\tau},\bar{\bm{\Gamma}},\nu)$, whereas the likelihood and posterior are $({\bf{y}}\mid{\bm{\beta}})\sim{{\mbox{SUT}}}_{n,q}({\bm{\xi}}_{{\bf{y}}\mid{\bm{\beta}}},\alpha_{\bm{\beta}}\bm{\Omega}_{{\bf{y}}\mid{\bm{\beta}}},{\bm{\Delta}}_{{\bf{y}}\mid{\bm{\beta}}},\alpha_{\bm{\beta}}^{-1/2}\bm{\tau}_{{\bf{y}}\mid{\bm{\beta}}},{\bar{\bm{\Gamma}}}_{{\bf{y}}\mid{\bm{\beta}}},\nu+p)$ and $({\bm{\beta}}\mid{\bf{y}})\sim{{\mbox{SUT}}}_{p,q}({\bm{\xi}}_{{\bm{\beta}}\mid{\bf{y}}},\alpha_{\bf{y}}\bm{\Omega}_{{\bm{\beta}}\mid{\bf{y}}},{\bm{\Delta}}_{{\bm{\beta}}\mid{\bf{y}}},\alpha_{\bf{y}}^{-1/2}\bm{\tau}_{{\bm{\beta}}\mid{\bf{y}}},{\bar{\bm{\Gamma}}}_{{\bm{\beta}}\mid{\bf{y}}},\nu+n)$, respectively. In these expressions, $\alpha_{\bm{\beta}}=[\nu+Q_{\bm{\beta}}({\bm{\beta}})]/(\nu+p)$ and $\alpha_{\bf{y}}=[\nu+Q_{\bf{y}}({\bf{y}})]/(\nu+n)$. The remaining parameters, along with $Q_{\bm{\beta}}({\bm{\beta}})$ and $Q_{\bf{y}}({\bf{y}})$, are defined as in Proposition 1.
@@
The proofs follow directly by replacing the generic density generators within Proposition 1 with those of the Student’s $t$ distribution presented in Section 2.2.2. Alternatively, it is possible to prove the statement leveraging the specific properties of unified skew-$t$ distributions in Proposition 11 of Wang et al. [62]. ∎
Corollary 1 states a general conjugacy result which further includes those of the SUN as a limiting case, provided that SUT distributions converge to SUNs when $\nu\to\infty$ [e.g., 62]. In addition, suitable constraints on the parameters of the joint SUT distribution for $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}$ in Corollary 1, yield priors and likelihoods of potential practical interest. In particular, consider $({\bm{\beta}}^{\top},{\bm{\varepsilon}}^{\top})^{\top}\sim{{\mbox{SUT}}}_{p+n,q}({\bm{\xi}},\bm{\Omega},{\bm{\Delta}},\bm{\tau},\bar{\bm{\Gamma}},\nu),$ with
@@
which yield a Bayesian multivariate regression model ${\bf{y}}={\bf{X}}{\bm{\beta}}+{\bm{\varepsilon}}$ with Student’s $t$ prior on ${\bm{\beta}}$ and unified skew-$t$ residuals in ${\bm{\varepsilon}}$, uncorrelated with ${\bm{\beta}}$. In the above expression, $\mathcal{T}_{p}({\bm{\xi}}_{\bm{\beta}},\bm{\Omega}_{\bm{\beta}},\nu)$ denotes the $p$-variate Student’s $t$ distribution with location ${\bm{\xi}}_{\bm{\beta}}$, scale $\bm{\Omega}_{\bm{\beta}}$ and degrees of freedom $\nu$. By Corollary 1, this implies the SUT likelihood